Shift Summary for EVENING shift 8/30/99

Hogan Nguyen (hogann@fnal.gov)
Tue, 31 Aug 1999 00:09:30 -0500 (CDT)

Shift Summary for DAY shift, 30/Aug/99

On Shift: Ricardo Barbosa, Hogan

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of E832 data (special runs not included): 8.0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Changing tapes and initializing a run;
drt_fresh_start; occasional daqdead warnings

========================================================================
Runs ended this shift:
Run 14378 - 19.4M events - 27.7 2pi0 - BEAM_832

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================
Accesses during shift:
None.

========================================================================
Overview of shift:
Steady beam, continuous data taking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14378 15:30 Inherit smooth E832 data taking.

15:55 Accelerator looses timing generator. No beam

16:00 Timing generator rebooted. Beam comes back.

16:25 Tapes full. Change tapes.

RSK requests to do drt_fresh_start inorder to
pick up new online code to deal with loss of
trigger latch crate. See log book XXI page 119.

14379 1 16:48 start BEAM_832 run

6 pmt gain drift looks weird for the first couple
of spills. Back to normal in spill 6.
JAWS says this is related to the ktevana errors
at the beginning of this run.

17:40 Paged Sydney to have him delete unneeded split
files. He calls back from his car while it is
in a car wash. We don't want him to have
run manager withdrawl.

20:02 Spills 146-148 lost due to lost permit in stream 5
Finally fixed by cycling repeater near by the
laser crate.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

o if we lose beam for >= 40 minutes, and the DA works, do a laser scan.

o keep your eyes on channel 450 (see log book p. 103). If it remains stable,
and we do a laser calibration, maybe we don't need to swap the divider
(ask Ed). If it is unstable, page Julie and Ed.

o trg#summary shows RC10 trigger source is bad. RSK says not to worry
since the extra triggers sources are o.k. Elizabeth has been notified.

o When tapes are full, do muon run and set NM4AN polarity to positive.